OpenStudy (anonymous):

an object falls a distance h from rest.if it travels 0.60h in the last 1.00s. find the time and the height of its fall.

OpenStudy (anonymous):

\[h = 9.81*t ^{2} = \frac{ 9.81 }{ 0.4 } * (t-1)^{2}\] \[t =\frac{\sqrt{1.6}}{1.2} and h = 9.81*t^{2}\]

OpenStudy (irishboy123):